OpenStudy (anonymous):

One number is nine more than a second number. Four times the first is 2 less than 6 times the second. Find the numbers. Am I right? (x+9)+y , (4x-2)+6y

OpenStudy (nastech):

in this case, you have to use a single variable.

OpenStudy (anonymous):

so how do i rewrite it then?

OpenStudy (nastech):

the first number is nine more than the second; (y+9) is the first number and then y is the second number

OpenStudy (nastech):

four times the first is two less than six times the second; {4(y+9)}=6y-2

OpenStudy (nastech):

do u get it

OpenStudy (anonymous):

i think so thank you so much!

OpenStudy (nastech):

what answer do u have

OpenStudy (anonymous):

im working it one second..

OpenStudy (anonymous):

19=y

OpenStudy (anonymous):

but i need 2 values for each equation so why did i only have on outcome?

OpenStudy (anonymous):

*number, not equation

OpenStudy (nastech):

4(y+9)=6y-2 4y+36=6y-2 10y=38 y=3.8 3.8 is the second number but we read that the first number is 9 more than the second. so to get the first number, add 9 to the second number which is 3.8

OpenStudy (nastech):

first number = 3.8+9=12.8 second number = 3.8

OpenStudy (anonymous):

why did you add 4y to 6y shouldnt you have subtracted?

OpenStudy (nastech):

sorry about that, i should have subtracted

OpenStudy (nastech):

first number= 19+9=28 second number=19

OpenStudy (nastech):

you were right

OpenStudy (anonymous):

awsome okay thank you for struggling through that with me lol:)

OpenStudy (nastech):

u are welcome